A sample of gas has a mass of 38.8 mg . Its volume is 228 mL at a temperature of 55 ∘C and a pressure of 829 torr . Find the molar mass of the gas.
Answer #1

Volume of gas V=228ml =0.228L

Temperature of gas T=55℃=55+273=328K

Pressure of gas P =829torr =829/760=1.09atm

(As 760torr=1atm)

By ideal gas law PV=nRT

Where n is the number of moles of gas

R=0.0820Latm/molK

n=PV/RT

Therefore n=(1.09 atm×0.228L)/(0.0820 Latm/molK ×328K)

=9.24×10-3mol

Given mass of gas=38.8mg= 38.8×10-3g

Given mass/molar mass = number of moles of gas

Molar mass =given mass/number of moles

=38.8×10-3/9.24×10-3 =4.199g/mol

Thank you.